# Experimental Contexts Can Facilitate Robust Semantic Property Inference in Language Models, but Inconsistently

### v1

@@ Abstract
[…]
@@ 1 Introduction
Carefully controlled behavioral analyses on meaning-sensitive tasks have revealed holes in the ability of language models (LMs) to demonstrate robust meaning extraction and use (Pandia and Ettinger, 2021; Elazar et al., 2021; Schuster and Linzen, 2022; Misra et al., 2023; Kim and Schuster, 2023, i.a). However, since a large subset of these investigations uses zero-shot evaluation as the primary methodology, there are growing concerns that they do not paint a complete picture of LMs’ abilities (Lampinen, 2022; Sinclair et al., 2022; Sinha et al., 2023). Conclusions that LMs lack a particular ability may be overhasty if it turns out the ability is easily accessed through in-context learning, different question formulations, or particular instructions (Lampinen, 2022; Wei et al., 2022).
Figure 1: We prompt LMs with in-context examples that are compatible with both, robust property inheritance, as well as position-based heuristics. At test time, we evaluate the model on cases where the heuristics support desirable behavior and on cases where they do not. We use stimuli from comps and its reformulation as a QA task (comps-qa).
Our focus [^1] Our code can be found in this link in this paper is a particularly challenging data set for meaning-sensitive behavior: comps (Misra et al., 2023), a dataset of minimal pair sentences that tests the ability of LMs on property knowledge of everyday concepts (a beaver/gorilla has a flat tail) and their inheritance for novel concepts (a wug is a beaver/gorilla. therefore a wug has a flat tail). Contemporary LMs failed miserably on the hardest subset of the comps stimuli, the examples of which contain two novel concepts (wug vs. dax), where only one of them inherits the target property (has a flat tail):
@@
Given the success of LMs on a wide variety of complicated tasks, their utter failure on this seemingly straightforward task remains puzzling. Here, we systematically explore comps on five modern LMs ranging from 1.5–13B parameters, varying (a) whether models are evaluated zero-shot or with multiple examples and (b) whether or not instructions are present.
Unlike other minimal-pair datasets, using comps in an in-context learning setting is non-trivial (and thus potentially informative). This is because the task can be solved using a position-based heuristic. For example, in one subset of comps, the target property is always attached to the first novel concept—like in figure 1. Importantly, models’ failures on comps were shown to be in part a result of models’ tendencies towards heuristic behavior: the performance of LMs is particularly bad when the distractor (a dax is a gorilla) is recent—i.e., autoregressive LMs show a recency bias in attributing properties to novel concepts. In that sense, comps follows a rich body of work in which tasks are set up in a manner that two types of generalization mechanisms can lead to the same prediction, but only one of which is desirable (McCoy et al., 2019, 2020; Warstadt et al., 2020b; Mueller et al., 2022; Si et al., 2023; Mueller et al., 2023).
We find that experimental contexts, as operationalized using in-context examples and instructions, can in fact demonstrate robust improvements in LMs’ property inheritance behavior as measured by the stimuli in comps. However, this improvement comes with a caveat: With a minimal reformulation of comps into a QA task, where there is a direct link between the LMs’ output space and the features of the input that control the heuristic, LMs show a strong preference towards the heuristic, and are therefore at chance. This discrepancy suggests that the improvements on the original task do not necessarily indicate that the models have successfully mastered the reasoning ability required to perform property inheritance, which remains a key challenge for them.
@@ 2 Methodology
We use the most difficult subset of the comps dataset (Misra et al., 2023)comps-wugs-dist—for our experiments. This dataset contains 13,828 sentence pairs of the form similar to figure 1, constructed using 152 animal concepts and 991 properties.
@@
We evaluated 5 different open-source LMs, all of which are decoder-only, and were accessed using the huggingface hub (Wolf et al., 2020): GPT-2 XL (Radford et al., 2019); OPT-6.7b (Zhang et al., 2022); Llama-2 (we used the 7b and the 13b versions; Touvron et al., 2023); and Mistral-7b (Jiang et al., 2023). Details about the models can be found in the appendix.
Figure 2: Overall results from our experiments testing LMs on comps and comps-qa using in-context examples, with and without instructions. Results are aggregated across both heuristics: first-correct and recent-correct. Error bars are over different sets of in-context examples. All models start off near chance in the 0-shot case, but many improve as more examples are given. Some (e.g., GPT-2 XL on comps-qa) show strongly heuristically driven behavior, as evidenced by the diverging performance on items where heuristics work and those where they do not. Figures 3 and 4 show fine-grained results.
@@ 3 Analyses and Results
On comps, models unsurprisingly start off at chance performance on average, corroborating the previous findings of Misra et al. (2023). However, in the presence of in-context examples and instructions, they are able to improve monotonically as the number of in-context examples increases. It is worth noting Llama-2-13b does occasionally show a slight preference for heuristics in the absence of instructions (e.g., 84% vs. 62% when prompted with 2 examples). An intermediate conclusion that we draw here is that LMs can indeed demonstrate non-trivial property inheritance on observing a few examples that reflect that behavior.
@@
While experimental context seems to aid models in attributing properties to the right concept in context, the same does not hold on comps-qa. Similar to comps, models start off at chance performance on average with a zero-shot set up, however, unlike in the case of comps, LMs seem to consistently prefer the heuristics available in the prompt, showing worse than chance performance on cases where the test set does not follow the heuristic. This is most apparent for GPT-2 XL, OPT6.7b, and Llama-2-7b—here the gap between the accuracy for cases where heuristics support performance on the test set and the accuracy for cases where they do not almost always worsens with an increase in the number of exemplars, on average. This is especially notable for OPT6.7b, which attains perfect performance on cases where the heuristics match up with the test set and at the same time it is close to 0% on cases where they do not. A notable exception to this trend is Mistral-7b, which seems to be resilient to the spurious heuristics, showing a net-positive improvement from the zero-shot case, especially in the presence of instructions. Nevertheless in the absence of instructions, it too shows a slight preference for position-based heuristics—for instance, its accuracy with 6 in-context exemplars when the heuristics support success on the test set is 82% and on cases where the heuristics oppose the test set is 65%.
@@
Our results suggest that LMs are more likely to show behavior that is compatible with the use of positional heuristics when their output space (choice between the two novel concepts) has a clear connection with positional artifacts in their input (relative ordering of the novel concepts). This is consistent with our hypothesis in about 8 out of 10 cases. When this link is not clear and models must instead predict likely properties given a novel concept (i.e., in comps), instructions and in-context examples do seem to lead to robust performance. It is important to note that instructions alone do not always account for the observed improvement—LMs’ performance on zero-shot settings are consistently still at chance in all cases, suggesting that it is the in-context examples that critically alter models’ output distribution to support desirable property inference behavior.
@@ 4 Conclusion
In this work, we investigated the extent to which in-context examples and instructions—key components that drive impressive performance in contemporary LMs—can overcome important limitations of LMs at tests that have poked holes in their ability to extract conceptual meaning from text. As a case study, we analyzed how well such experimental contexts can improve LM abilities to perform property inheritance (Murphy, 2002; Misra et al., 2023) in context—binding of novel concepts to existing concepts, and endowing them with valid property inferences as a result. Our findings suggest that mastery of this ability has yet to be robustly achieved, and that LMs in general are still prone to using shallower patterns in their context (when available) rather than systematically extracting conceptual meaning. At the same time, exploring precisely what makes Mistral less susceptible to heuristics will be useful to design more robust LMs, which we leave for future work.
@@ 5 Limitations
A clear limitation of this work is that it exclusively focuses on a single dataset: comps (Misra et al., 2023). So, a question that arises here is to what extent are our findings localized to the chosen dataset vs. meaning-sensitive evaluations in general. This would require a further non-trivial, non-straightforward amount of work, since: (1) different meaning sensitive evaluations focus on different (though equally useful) operationalizations of meaning; and more importantly (2) not all prior work in this area focuses on a standardized and well-defined usage of heuristics that is directly transferable to the experimental setup we have used in this work (following McCoy et al., 2019, 2020; Warstadt et al., 2020b; Mueller et al., 2022; Si et al., 2023).
We do hope that our work contributes to the larger-scale vision of carefully benchmarking different types of meaning extraction abilities in LMs in a controlled manner.
@@ 6 Acknowledgments
We thank Tom McCoy and Andrew Lampinen for providing comments on a previous version of the draft. (KM)${}^{2}$ acknowledge funding from NSF Grant 2104995 awarded to Kyle Mahowald.
@@ A.1 Methodological details
[…]
@@
In both cases above, gek is the concept that should inherit the property. While these examples show the zero-shot case, cases with in-context examples and instructions simply add more context to the prefix, therefore the surface form of the output space remains the same regardless of the number of in-context examples or the presence of instructions.
@@ A.2 Model Metadata
GPT-2 XL 1.5B 8B 50,257
OPT-6.7b 6.7B 180B 50,272
Llama-2-7b 7B 2T 32,000
Llama-2-13b 13B 2T 32,000
Mistral-7b 7B ? 32,000
Table 1: Overview of the LMs used in this work. ‘?’ indicates that the given value was not made available in the LM’s release.
@@ Appendix B Instructions
comps
Given a pair of statements that introduce novel entities as types of real world animals, write a true statement about the properties of the novel entities:
{exemplars} (omitted in zero-shot)
@@
comps-qa
Given a pair of statements that introduce novel entities as types of real world animals, answer the question that follows:
{exemplars} (omitted in zero-shot)
@@
comps
Some aliens have come to earth, and it turns out they have their own language for talking about our animals here on Earth. Your job is to help the aliens learn about our Earthling animals by giving them some information about the animals.
@@
{exemplars} (omitted in zero-shot)
@@
comps-qa
Some aliens have come to earth, and it turns out they have their own language for talking about our animals here on Earth. Your job is to help the aliens learn about our Earthling animals by answering some questions about them.
@@
{exemplars} (omitted in zero-shot)
@@
comps Zero-shot
It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world animals. Your task is to make a conclusion about the properties of one of the entities by reasoning over the premise statements.
@@
Few-shot
It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals, followed by another statement that attributes a property to one of the entities introduced in the premise statements.
@@
comps-qa Zero-shot
It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world objects. The statements are followed by a question that asks which novel entity in the premise can a specific property can be attributed to. Answer the question by reasoning over the premise statements.
@@
Few-shot
It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals. The statements are followed by a question that asks which novel entity in the premise can a specific property can be attributed to, and the answer to the question, obtained by reasoning over the premise statements.
@@
comps Zero-shot
It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world animals. Your task is to write a true statement about the properties of the novel entities.
@@
Few-shot
It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals, followed by another statement that attributes a property to one of the entities introduced in the premise statements.
@@
comps-qa Zero-shot
It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world animals. The statements are followed by a question that asks which of the introduced entities a specific property can be attributed to. Answer the question by reasoning over the premise statements.
@@
Few-shot
It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals. The statements are followed by a question that asks which of the introduced entities a specific property can be attributed to, and the answer to the question, obtained by reasoning over the premise statements.

### v2

@@ Abstract
Kanishka Misraτ  Allyson Ettingerα  Kyle Mahowaldτ τThe University of Texas at Austin  αAllen Institute for Artificial Intelligence kmisra@utexas.eduallysone@allenai.orgkyle@utexas.edu
@@ 1 Introduction
Carefully controlled behavioral analyses on meaning-sensitive tasks have revealed holes in the ability of language models (LMs) to demonstrate robust meaning extraction and use (Pandia and Ettinger, 2021; Elazar et al., 2021; Schuster and Linzen, 2022; Weissweiler et al., 2022; Misra et al., 2023; Kim and Schuster, 2023, i.a). However, since a large subset of these investigations uses zero-shot evaluation as the primary methodology, there are growing concerns that they do not paint a complete picture of LMs’ abilities (Lampinen, 2022; Sinha et al., 2023). Conclusions that LMs lack a particular ability may be overhasty if it turns out the ability is easily accessed through in-context learning, different question formulations, or particular instructions (Lampinen, 2022; Wei et al., 2022).
Figure 1: LMs are prompted with in-context examples that are compatible with both, robust property inheritance, as well as position-based heuristics. At test time, we evaluate on cases where the heuristics support desirable behavior and on cases where they do not. We use stimuli from comps and its reformulation as a QA task.
Our focus [^1] Our code can be found in this repository in this paper is a particularly challenging dataset for meaning-sensitive behavior: comps (Misra et al., 2023), which contains minimal pair sentences that test the ability of LMs on property knowledge of everyday concepts (a beaver/gorilla has a flat tail) and their inheritance for novel concepts (a wug is a beaver/gorilla. therefore, a wug has a flat tail). Contemporary LMs failed miserably on the hardest subset of the comps stimuli, the examples of which contain two novel concepts (wug vs. dax), where only one of them inherits the target property (has a flat tail):
@@
Given the success of LMs on a wide variety of complicated tasks, their utter failure on this seemingly straightforward task remains puzzling. Here, we systematically explore comps on 12 LMs ranging from 1.5–13B parameters, varying (a) whether they are evaluated zero-shot or with multiple examples, and (b) whether or not instructions are present.
Unlike other minimal-pair datasets, using comps in an in-context learning setting is non-trivial (and thus potentially informative). This is because the task can be solved using a position-based heuristic. For example, in one subset of comps, the target property is always attached to first novel concept—like in figure 1. Importantly, LMs’ failures on comps were shown to be in part a result of models’ tendencies towards heuristic behavior: the performance of autoregressive LMs was found to be particularly bad when the distractor (a dax is a gorilla) is recent—i.e., they showed a recency bias in attributing properties to novel concepts. In that sense, comps follows a rich body of work in which tasks are set up in a manner that two types of generalization mechanisms can lead to the same prediction, but only one of which is desirable (McCoy et al., 2019, 2020; Warstadt et al., 2020b; Mueller et al., 2022; Si et al., 2023).
We find that experimental contexts, as operationalized using in-context examples and instructions, can in fact demonstrate robust improvements in LMs’ property inheritance behavior as measured by comps. However, this improvement comes with a caveat: With a minimal reformulation of comps into a QA task, where there is a direct link between the LMs’ output space and the features of the input that control the heuristic, many LMs showed a strong preference towards the heuristic, and were therefore at chance. Interestingly, LMs that receive explicit supervision to follow instructions do show some resistance to positional heuristics, but not always—occasionally, certain instruction-tuned LMs also tend to show a preference for one particular type of heuristic. These discrepancies in LMs’ performance underscore their difficulty of mastering the reasoning ability to robustly demonstrate semantic property inheritance.
@@ 2 Methodology
We use the most difficult subset of the comps dataset (Misra et al., 2023)comps-wugs-dist. This dataset contains 13,828 sentence pairs of the form similar to figure 1, constructed using 152 animal concepts and 991 properties.
@@
We evaluated 8 different open-source LMs, all of which are decoder-only, and were accessed using the huggingface hub (Wolf et al., 2020): GPT-2 XL (Radford et al., 2019); OPT-6.7b (Zhang et al., 2022); Llama-2 (we used the 7B and the 13B versions; Touvron et al., 2023); Mistral-7B (Jiang et al., 2023); Llama-3-8B (Dubey et al., 2024); OLMo-7B (Groeneveld et al., 2024); and Gemma-2-9B (Team et al., 2024). In addition to these 8, we tested the instruct-tuned versions of OLMo, Llama-3-8B, Mistral-7B (v0.3), and Gemma-2-9B, to analyze if instruction tuning affects LMs’ robustness. Details about all 12 LMs can be found in Appendix A.
Figure 2: Overall results from our experiments testing non-instruction tuned LMs on comps and comps-qa using in-context examples, with and without instructions. Results are aggregated across both heuristics: first-correct and recent-correct. Error bars are over different sets of in-context examples. Most models start off near chance in the 0-shot case, but many improve as more examples are given. Solid green line depicts each model’s base property knowledge performance, while the black dashed line depicts chance performance.
@@ 3 Analyses and Results
On comps, models unsurprisingly start off at chance performance on average, corroborating the previous findings of Misra et al. (2023). However, in the presence of in-context examples and instructions, they are able to improve monotonically as the number of in-context examples increases, reaching similar levels of performance on property inheritance as that of the models’ base knowledge. It is worth noting Llama-2-13b and OLMo-7B do occasionally show a preference for heuristics in the absence of instructions. An intermediate conclusion that we draw here is that LMs can indeed demonstrate non-trivial property inheritance on observing a few examples that reflect that behavior.
@@
While experimental context seems to aid models in attributing properties to the right concept in context, the same does not hold on comps-qa. Similar to comps, models start off at chance performance on average with a zero-shot set up, however, unlike in the case of comps, many LMs seem to consistently prefer the heuristics available in the prompt, showing, at times, worse than chance performance on cases where the test set does not follow the heuristic. This is most apparent for GPT-2 XL, OPT6.7b, Llama-2-7b, OLMo-7B—here there is a substantial gap between the accuracy for cases where heuristics support performance on the test set and the accuracy for cases where they do not. Other models, like Mistral-7B and Gemma-2-9B occasionally show behavior compatible with the heuristics, but this does not hurt their overall performance, which is consistently above chance, and close to their base knowledge.
@@
Our results suggest that LMs are more likely to show behavior that is compatible with the use of positional heuristics when their output space (choice between the two novel concepts) has a clear connection with positional artifacts in their input (relative ordering of the novel concepts). This is consistent with our hypothesis in about 11 out of 16 cases (where for 9 of the cases, heuristics end up drastically hurting performance). When this link is not clear and models must instead predict likely properties given a novel concept (i.e., in comps), instructions and in-context examples do seem to lead to robust performance. It is important to note that instructions alone do not always account for the observed improvement—LMs’ performance on zero-shot settings are consistently still at chance in all cases, suggesting that it is their interaction with in-context examples that critically alters models’ output distribution to support desirable property inference behavior.
Figure 3: Overall results on the four instruction-tuned models considered. Results are aggregated across both heuristics: first-correct and recent-correct. Error bars are over different sets of in-context examples. Solid green line depicts each model’s base property knowledge performance, while the black dashed line depicts chance performance.
Benefits of instruction tuning…
Common to the previous analysis is that the tested LMs were all pretrained using the standard next-word prediction objective. While instruction-following capabilities have been shown to arise in such models, a new paradigm has emerged in the field that tunes such pre-trained LMs post-hoc to explicitly follow instructions using a plethora of different techniques not limited to those that leverage feedback from human-provided labels (Ouyang et al., 2022). How does instruction tuning interact with the presence of position-based heuristics? Figure 3 shows results on four LMs which are instruction-tuned versions of some of the LMs used in the previous analysis, for comps and comps-qa (with instructions). We find these four LMs to show consistently above-chance performance on both datasets, suggesting non-trivial robustness brought about by instruction-tuning. Interestingly, while most models show improvements from zero-shot (often times at chance), the instruct-tuned Gemma-2-9B model shows no signs of improvement, instead showing nearly constant performance on both comps and comps-qa regardless of the number of in-context examples.
…and its occasional pitfalls
While Figure 3 shows that LMs benefit from instruction-tuning in the context of robust property inheritance, a detailed breakdown of the results offers a few caveats. In Figure 4 we show OLMo-7B-Instruct’s accuracies broken-down by heuristic on comps-qa, the setting where we hypothesized LMs to be most likely relying on heuristics. For the first-correct heuristic, the model shows a preference for the heuristic (greater performance when it supports success on the test set), while for recent-correct, we see the opposite to be true. Here, we observe that the LM does better when the heuristic does not support success on the test set, suggesting that the LM has a bias for preferring entities mentioned first—a bias so strong that it persists even when there is ample evidence in the context of the model that favors the opposite heuristic. This suggests that while instruction tuning leads to consistently above-chance performance on challenging property inheritance problems, it is not entirely robust to position-based heuristics.
Figure 4: Finer-grained results on Instruct-tuned OLMo-7B LM demonstrating its preference for selecting the first concept, regardless of the heuristics.
@@ 4 Conclusion
We investigated the extent to which in-context examples and instructions—key components that drive impressive performance in contemporary LMs—can overcome important limitations of LMs at tests that have poked holes in their ability to extract conceptual meaning from text. As a case study, we analyzed how well such experimental contexts can improve LM abilities to perform property inheritance (Murphy, 2002; Misra et al., 2023) in context: binding of novel concepts to existing concepts, and endowing them with valid property inferences as a result.
Our findings suggest that mastery of this ability has yet to be robustly achieved, and that many LMs trained on the next-word prediction objective are still prone to using shallower patterns in their context rather than systematically extracting conceptual meaning. There are several exceptions to this, however, for example, modern LMs such as Mistral-7B and Gemma-2-9b did show heuristic behavior, but were also well above chance, especially in the presence of instructions. Additionally, instruction tuned LMs were found to be substantially more robust than their non-instruction tuned counterparts, suggesting advantages to explicit instruction tuning. However, they too sometimes showed counterintuitive behavior, where their strong bias for a consistent position overcomes the heuristic patterns available in the input. We leave the nuanced exploration of this behavior induced by instruction-tuning for future work.
@@ 5 Limitations
A clear limitation of this work is that it exclusively focuses on a single dataset: comps (Misra et al., 2023). So, a question that arises here is to what extent are our findings localized to the chosen dataset vs. meaning-sensitive evaluations in general. This would require a further non-trivial, non-straightforward amount of work, since: 1) different meaning sensitive evaluations focus on different (though equally useful) operationalizations of meaning; and more importantly 2) not all prior work in this area focuses on a standardized and well-defined usage of heuristics that is directly transferable to the experimental setup we have used in this work (following McCoy et al., 2019, 2020; Warstadt et al., 2020b; Mueller et al., 2022; Si et al., 2023). We do hope that our work contributes to the larger-scale vision of carefully benchmarking different types of meaning extraction abilities in LMs in a controlled manner.
@@ 6 Acknowledgments
We thank Tom McCoy and Andrew Lampinen for providing comments on a previous version of the draft. (KM)2 acknowledge funding from NSF Grant 2104995 awarded to Kyle Mahowald. We thank UT-Austin Linguistics for providing the hardware to run experiments.
@@ A.1 Methodological details
$\log P_{\theta}(\textit{{{\color[rgb]{0.23828125,0.5234375,0.77734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.23828125,0.5234375,0.77734375}gek}/{\color[rgb]{0.9453125,0.76171875,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9453125,0.76171875,0.1953125}wif}}}\mid$ a gek is a beaver. a wif is a gorilla.
@@
In both cases above, gek is the concept that should inherit the property. While these examples show the zero-shot case, cases with in-context examples and instructions simply add more context to the prefix, therefore the surface form of the output space remains the same regardless of the number of in-context examples or the presence of instructions. Similar to the above measures, we measure the base property knowledge of LMs by comparing the conditional log-probabilities of the property phrases given the real, known concepts in question. That is, we measure:
[…]
The accuracy here, again, is the proportion of time the correct concept was more likely to inherit the property than the incorrect one. We measure this accuracy for all our test items zero-shot, and use that as the reference base property knowledge in order to contextualize our main results.
@@ A.2 Model Metadata
GPT-2 XL (Radford et al., 2019) 1.5B 8B 50,257
OPT-6.7b (Zhang et al., 2022) 6.7B 180B 50,272
Llama-2-7B (Touvron et al., 2023) 7B 1.8T 32,000
Mistral-7B (Jiang et al., 2023) 7B ? 32,000
OLMo-7B (Groeneveld et al., 2024) 7B 2.46T 50,304
Llama-3-8B (Dubey et al., 2024) 8B 15T 128,000
Gemma-2-9B (Team et al., 2024) 9B 8T 256,000
Llama-2-13B (Touvron et al., 2023) 13B 1.8T 32,000
Table 1: Overview of the non-instruction tuned LMs used in this work. ‘?’ indicates that the given value was not made available in the LM’s release. In addition to these, we also report results on instruction-tuned versions of OLMo-7B, Mistral-7B, Llama-3-8B, and Gemma-2-9B.
@@ Appendix B Instructions
comps Given a pair of statements that introduce novel entities as types of real world animals, write a true statement about the properties of the novel entities:
{examples} (omitted in zero-shot)
@@
comps-qa Given a pair of statements that introduce novel entities as types of real world animals, answer the question that follows:
{examples} (omitted in zero-shot)
@@
comps Some aliens have come to earth, and it turns out they have their own language for talking about our animals here on Earth. Your job is to help the aliens learn about our Earthling animals by giving them some information about the animals.
@@
{examples} (omitted in zero-shot)
@@
comps-qa Some aliens have come to earth, and it turns out they have their own language for talking about our animals here on Earth. Your job is to help the aliens learn about our Earthling animals by answering some questions about them.
@@
{examples} (omitted in zero-shot)
@@
comps Zero-shot It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world animals. Your task is to make a conclusion about the properties of one of the entities by reasoning over the premise statements.
@@
Few-shot It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals, followed by another statement that attributes a property to one of the entities introduced in the premise statements.
@@
comps-qa Zero-shot It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world objects. The statements are followed by a question that asks which novel entity in the premise can a specific property can be attributed to. Answer the question by reasoning over the premise statements.
@@
Few-shot It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals. The statements are followed by a question that asks which novel entity in the premise can a specific property can be attributed to, and the answer to the question, obtained by reasoning over the premise statements.
@@
comps Zero-shot It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world animals. Your task is to write a true statement about the properties of the novel entities.
@@
Few-shot It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals, followed by another statement that attributes a property to one of the entities introduced in the premise statements.
@@
comps-qa Zero-shot It is important to know and reason about the properties of entities in the world. The following is a pair of premise statements that introduce novel entities as new types of real world animals. The statements are followed by a question that asks which of the introduced entities a specific property can be attributed to. Answer the question by reasoning over the premise statements.
@@
Few-shot It is important to know and reason about the properties of entities in the world. The following example(s) show a pair of premise statements that introduce novel entities as new types of real world animals. The statements are followed by a question that asks which of the introduced entities a specific property can be attributed to, and the answer to the question, obtained by reasoning over the premise statements.